If angle 1 and angle 2 are supplementary angles and angle 1 is 12 degrees more than angle 2, what is the measurement of each angle?
Answer by stanbon(75887)

If angle 1 and angle 2 are supplementary angles and angle 1 is 12 degrees more than angle 2, what is the measurement of each angle?
------------------
x + y = 180
x = y+12
------
Modify::
x + y = 180
x-y = 12
----
Add and solve for "x":
2x = 192
x = 96 degrees
-----
y = 84 degrees
===============
Cheers,
Stan H.
===============